# Schrödinger cats escape confinement

### v2

@@ Abstract
We consider local quenches from initial states generated by a single spin-flip in either the true or the false vacuum state of the confining quantum Ising spin chain in the ferromagnetic regime. Contrary to global quenches, where the light-cone behaviour is strongly suppressed, we find a significant light-cone signal propagating with a nonzero velocity besides the expected localised oscillating component. Combining an analytic representation of the initial state with a numerical description of the relevant excitations using the two-fermion approximation, we can construct the spectrum of post-quench excitations and their overlaps with the initial state, identifying the underlying mechanism. For confining quenches built upon the true vacuum, the propagating signal consists of Schrödinger cats of left and right-moving mesons escaping confinement. In contrast, for anti-confining quenches built upon the false vacuum, it is composed of Schrödinger cats of left and right-moving bubbles which escape Wannier-Stark localisation.
@@ 1 Introduction
Here, we study the dynamics of the ferromagnetic Ising spin chain with local quenches starting from initial states obtained by flipping a single spin. An advantage of these protocols is that dynamical confinement can be entirely separated from Wannier-Stark localisation, the former appearing in confining quenches where the initial magnetisation is aligned with the post-quench longitudinal field, while the latter arising when they are opposite. These correspond to the initial states obtained by a local change in the post-quench Hamiltonian’s true/false vacuum state. In contrast to quenches starting from a domain wall, we observe very prominent escaping fronts tracing out a well-defined light cone in the time evolution of the magnetisation. We explain the origin of these escaping fronts using a combination of exact analytic and numerical approaches. This leads to the conclusion that the relevant states are equal weight superposition of left and right-moving excitations, i.e., Schrödinger cats.
@@ 2 Local quenches in the confining Ising spin chain
in the two ground states $\ket{\pm}$ which are degenerate in the thermodynamic limit.
@@ 2.1 Confinement in the Ising spin chain
Choosing for definiteness $m>0$, the true ground state of the system is a perturbation of the one with positive magnetisation $\ket{+}$, while $\ket{-}$ becomes a metastable state, a.k.a. the false vacuum. In the following, we stay within the scenario of strong confinement, which corresponds to a transverse field $h_{t}$ far below its critical value $1$ and consider values of the longitudinal field $h_{l}$ of order $0.1$.
@@
To realise a confining quench, one starts from a ground state polarised parallel to the longitudinal field in the post-quench Hamiltonian. As shown in [3], confining forces radically alter the time evolution after a global quantum quench, suppressing the characteristic light-cone spreading of correlations. The resulting suppression is incomplete: string breaking leads to multi-meson excitations escaping confinement. However, this effect is usually strongly suppressed: in the quenches within the ferromagnetic case considered in the original work [3], the “escaping fronts” observed in the two-point correlation functions were suppressed by several orders of magnitude compared to the localised component. The reason behind this suppression is well-understood [55, 27, 56]: string breaking is analogous to pair creation in a homogeneous electric field known as the Schwinger effect [57], with the role of the electric field played by the string tension (2.4). For small values of $h_{t}$ and $h_{l}$, the rate of string breaking is approximately given by [26]
@@ 2.2 Phenomenology of local quenches
[…]
@@
where $k_{0}>0$ gives the initial value of the kink momenta as $k_{1,2}(t=0)=\pm k_{0}$, while $d_{0}>0$ gives their initial position as $x_{1,2}(t=0)=\pm d_{0}+X$, with their center located at $X$. In the local quenches considered here, the typical value of $d_{0}$ is expected to be of the order of the lattice spacing. At the same time, the enveloping curve of the localised component is given by the trajectory with $k_{0}=\pi$. Fig. 2.1 shows the predicted enveloping curve with red lines choosing $d_{0}=1$.
@@
Figure 2.3: The front propagation in the longitudinal magnetisation $\langle\hat{\sigma}_{j}^{x}\rangle$ for $h_{t}=0.25$ and $h_{l}\in\{-0.1,\,-0.2,\,-0.3,\,-0.4\}$ respectively, starting from the initial state |$\ket{\overline{\Psi}(0)}=\sigma^{z}_{L/2}\ket{+}_{h_{t},h_{l}=0}$. The solid lines indicate the maximum bubble velocities calculated from the corresponding bubble spectrum as explained in Subsection 3.2, while the dashed lines indicate the maximum free kink velocity.
Figure 2.4: The front propagation in the magnetisation for $h_{t}=0.25$ and $h_{l}\in\{-0.1,\,-0.2,\,-0.3,\,-0.4\}$ respectively, starting from the initial state |$\ket{\overline{\Psi}(0)}=\sigma^{z}_{L/2}\ket{+}_{h_{t},h_{l}=0}$. The red curves indicate the corresponding quasi-classical trajectories of the kink pair with initial momentum $k_{0}=0$.
@@
Despite the repulsive force between the kinks characteristic for this case, Bloch oscillations lead to Wannier-Stark localisation of the dynamics [26, 38], with the resulting evolution displayed in Figs. 2.3 and 2.4. While it looks superficially similar to the case of confining quenches illustrated in Fig. 2.1, we reveal significant differences in the underlying mechanism behind the escaping fronts.
@@ 3.1 Confinement: mesons
It is clear from Fig. 3.1 that these functions have a maximum value inside the Brillouin zone. The existence of this maximum is guaranteed by the Lieb-Robinson theorem [14], and its value depends on the species $n$. For collisionless mesons (large $n$), the independence of their energy from their momentum $K$ implies that these mesons are stationary irrespective of the value of $K$, which is nicely explained by the semi-classical description. In addition, the velocities monotonously decrease with $n$; hence, the light-cone bounding the escaping fronts in Fig. 2.1 corresponds to the maximum velocity of the lightest meson species $M_{1}$.
@@ 3.2 Anti-confinement: bubbles
Note that in contrast to the mesons, the sign of the velocity is opposite to that of the momentum. However, these velocities still have finite maxima; the largest of these maximum velocities corresponds to the first bubble $B_{1}$, which is the one corresponding to the light-cone boundary of anti-confining quenches in Fig. 2.3.
@@ 4.1 Initial state for $h_{l}=0$
$|+\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle_{\mathrm{NS}}+|0\rangle_{\mathrm{R}}\right)\,.$ (4.1)
@@
$|\Psi(0)\rangle=\hat{\sigma}^{z}_{L/2}|+\rangle=\frac{1}{\sqrt{2}}\left(|\Psi(0)\rangle_{\mathrm{NS}}+|\Psi(0)\rangle_{\mathrm{R}}\right)\,,$ (4.2)
@@
$\hat{\sigma}^{z}_{L/2}|-\rangle=\frac{1}{\sqrt{2}}\left(|\Psi(0)\rangle_{\mathrm{NS}}-|\Psi(0)\rangle_{\mathrm{R}}\right)\,,$ (4.14)
@@
$|-\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle_{\mathrm{NS}}-|0\rangle_{\mathrm{R}}\right)\,.$ (4.15)
@@ 4.2 Matching the meson wave functions
In this formula, $\beta_{n}(K)_{L}$ is the expected, non-physical overall phase that needs to be fixed for every state separately and depends on accidental phase choices for the eigenvectors obtained using numerical diagonalisation. [^4] While the knowledge of these phases is not essential in our considerations, they must be taken into account if one wishes to reproduce the time evolutions displayed in Figs. 2.1 and 2.3 in terms of the overlaps. However, there is a non-trivial momentum-dependent relative phase factor in (4.17), which can be attributed to a relative phase between the Bogolyubov fermions $\eta_{q}$ definition on the chain and those in the quantum mechanical description. Denoting the latter by $a_{q}$, the matching rules (4.19) correspond to the relation
@@ 4.3 Meson overlaps
which has a small global component. An example of such a quench is presented in Fig. C.1 a) of Appendix C, for the parameters $h_{t}=0.25$ and $h_{l}=0.2$. To the naked eye, there is no obvious difference between the time evolutions of the longitudinal magnetisation shown in Fig. C.1 a) and of the purely local quench shown in the second panel of Fig. 2.1. Indeed, the deviation turns out to be suppressed by around $3$ orders of magnitude compared to the signal from the localised component of the purely local quench.
@@
$\mathcal{C}_{n}(K)=\frac{1}{\sqrt{L}}\sum\limits_{\begin{subarray}{c}k\,\in\mathrm{NS/R}\\ k\neq 0,\pm K/2\end{subarray}}\sigma_{K/2-k}\sigma_{K/2+k}{\psi}_{n,K}^{*}(k)_{L}$ ${\mathcal{K}}(K/2-k,K/2+k)_{L}$ (4.30)
@@
Figure 4.1: Comparison of the meson overlaps with the (flipped) true vacuum state $\sigma^{z}_{L/2}\ket{+}_{h_{t},h_{l}=0}$ calculated applying exact diagonalisation (shown by crosses), and the meson overlaps with the flipped true vacuum state of the purely transverse system $\hat{\sigma}^{z}_{L/2}|+\rangle$ (given by (4.2)) calculated via the semi-analytic method (shown by lines) for the cases $h_{t}=0.25$, $h_{l}=0.1$ and $h_{l}=0.3$.
@@
The first two terms describe zero-momentum mesons building up the localised oscillations visible in Fig. 2.2. At the same time, the sum contains single mesons in a superposition of a left-moving and a right-moving state with the same velocity contributing with equal weights, corresponding to Schrödinger cats escaping confinement.
Figure 4.2: Overlaps of meson states |$\ket{M_{n}(K)}$ with the state $\hat{\sigma}^{z}_{L/2}|+\rangle$ (given by (4.2)) as a function of $n$ at different values of momentum $K$.
@@ 5 Escaping fronts in anti-confining local quenches
$\overline{\mathcal{C}}_{n}(K)$ $=\frac{1}{\sqrt{L}}\sum\limits_{\begin{subarray}{c}k\,\in\mathrm{NS}\end{subarray}}\sigma_{K/2-k}\sigma_{K/2+k}\overline{{\psi}}_{n,K}^{*}(k)_{L}{\mathcal{K}}(K/2-k,K/2+k)_{L}$ (5.7)
@@
Figure 5.1: Comparison of the bubble overlaps with the (flipped) false vacuum state of the pure transverse system $\sigma^{z}_{L/2}\ket{-}_{h_{t},h_{l}=0}$ calculated from exact diagonalisation (crosses) with the bubble overlaps with the state $\sigma^{z}_{L/2}\ket{-}$ (given by (4.14)) calculated via the semi-analytic method, at the parameter values $h_{t}=0.25$ and $h_{l}=0.1$ (continuous lines).
@@
Figure 5.2: Overlaps of bubble states |$\ket{B_{n}(K)}$ with the with the state $\sigma^{z}_{L/2}\ket{-}$ (given by (4.14)) as a function of $n$ at different values of momentum $K$.
@@
The latter conclusion is further supported by the fact that the maximum velocity of the smallest bubble $B_{1}$ matches the escaping fronts well, as shown in Fig. 2.3.
To sum up, for the anti-confining quenches, dynamical confinement is replaced by Wannier-Stark localisation, but again, it is the Schrödinger cats which escape it, with the only difference that these are bubble cats instead of mesonic ones.
@@ 6 Conclusions and outlook
In contrast with global quenches, our simulations show that after local quenches obtained by flipping a single spin in the ground state, the magnetisation has a significant propagating component besides the localised one. To determine the origin of these fronts, we computed the overlaps of meson states with the initial state using a semi-analytic approach, which works for any chain length, including the thermodynamic limit. The conclusion is that the light-cone fronts originate from states containing a single meson excitation in superpositions of left and right-moving mesons (mainly the lightest one) with equal weights, which is in contrast to global quenches where quasi-particle pairs trace out the light cones and are strongly suppressed by confinement.

### v3

# Escaping fronts in local quenches of a confining spin chain

@@ Abstract
We consider local quenches from initial states generated by a single spin-flip in either the true or the false vacuum state of the confining quantum Ising spin chain in the ferromagnetic regime. Contrary to global quenches, where the light-cone behaviour is strongly suppressed, we find a significant light-cone signal propagating with a nonzero velocity besides the expected localised oscillating component. Combining an analytic representation of the initial state with a numerical description of the relevant excitations using the two-fermion approximation, we can construct the spectrum of post-quench excitations and their overlaps with the initial state, identifying the underlying mechanism. For confining quenches built upon the true vacuum, the propagating signal consists of superpositions of left and right-moving mesons escaping confinement. In contrast, for anti-confining quenches built upon the false vacuum it is composed of superpositions of left and right-moving bubbles which escape Wannier-Stark localisation.
@@ 1 Introduction
Here, we study the dynamics of the ferromagnetic Ising spin chain with local quenches starting from initial states obtained by flipping a single spin. An advantage of these protocols is that dynamical confinement can be entirely separated from Wannier-Stark localisation, the former appearing in confining quenches where the initial magnetisation is aligned with the post-quench longitudinal field, while the latter arising when they are opposite. These correspond to the initial states obtained by a local change in the post-quench Hamiltonian’s true/false vacuum state. In contrast to quenches starting from a domain wall, we observe very prominent escaping fronts tracing out a well-defined light cone in the time evolution of the magnetisation. We explain the origin of these escaping fronts using a combination of exact analytic and numerical approaches. This leads to the conclusion that the relevant states are equal-weight superpositions of left and right-moving excitations.
@@ 2 Local quenches in the confining Ising spin chain
in the two ground states |±$\ket{\pm}_{h_{t},h_{l}=0}$ which are degenerate in the thermodynamic limit.
@@ 2.1 Confinement in the Ising spin chain
Choosing for definiteness $m>0$, the true ground state of the system is a perturbation of the one with positive magnetisation |$\ket{+}_{h_{t},h_{l}=0}$, while |$\ket{-}_{h_{t},h_{l}=0}$ becomes a metastable state, a.k.a. the false vacuum. In the following, we stay within the scenario of strong confinement, which corresponds to a transverse field $h_{t}$ far below its critical value $1$ and consider values of the longitudinal field $h_{l}$ of order $0.1$.
@@
To realise a confining quench, one starts from a ground state polarised parallel to the longitudinal field in the post-quench Hamiltonian. As shown in [3], confining forces radically alter the time evolution after a global quantum quench, suppressing the characteristic light-cone spreading of correlations. Note that due to the translational invariance of the initial state after a global quench, only meson states with zero total momentum eigenvalue can contribute; therefore, the one-meson states are localised. Nevertheless, string breaking leads to multi-meson excitations (with zero total momentum), producing fronts escaping confinement. However, this effect is usually strongly suppressed: in the quenches within the ferromagnetic case considered in the original work [3], the “escaping fronts” observed in the two-point correlation functions were smaller by several orders of magnitude compared to the localised component. The reason behind this suppression is well-understood [55, 27, 56]: string breaking is analogous to pair creation in a homogeneous electric field known as the Schwinger effect [57], with the role of the electric field played by the string tension (2.4). For small values of $h_{t}$ and $h_{l}$, the rate of string breaking is approximately given by [26]
@@ 2.2 Phenomenology of local quenches
(a) The solid lines indicate the maximum meson velocities calculated from the corresponding meson spectrum as explained in Subsection 3.1, while the dashed lines indicate the maximum free kink velocity.
(b) The red curves indicate the corresponding quasi-classical trajectories of the quasi-particles with initial momenta $k_{0}=\pm\pi$.
@@
where $k_{0}>0$ gives the initial value of the kink momenta as $k_{1,2}(t=0)=\pm k_{0}$, while $d_{0}>0$ gives their initial position as $x_{1,2}(t=0)=\pm d_{0}+X$, with their center located at $X$. In the local quenches considered here, the typical value of $d_{0}$ is expected to be of the order of the lattice spacing. At the same time, the enveloping curve of the localised component is given by the trajectory with $k_{0}=\pi$. Fig. 1(b) shows the predicted enveloping curve with red lines choosing $d_{0}=1$.
@@
(a) The solid lines indicate the maximum bubble velocities calculated from the corresponding bubble spectrum as explained in Subsection 3.2, while the dashed lines indicate the maximum free kink velocity.
(b) The red curves indicate the corresponding quasi-classical trajectories of the kink pair with initial momentum $k_{0}=0$.
Figure 2.2: The front propagation in the longitudinal magnetisation $\langle\hat{\sigma}_{j}^{x}\rangle$ for $h_{t}=0.25$ and $h_{l}\in\{-0.1,\,-0.2,\,-0.3,\,-0.4\}$ respectively, starting from the initial state |$\ket{\overline{\Psi}(0)}=\sigma^{z}_{L/2}\ket{+}_{h_{t},h_{l}=0}$.
@@
Despite the repulsive force between the kinks characteristic for this case, Bloch oscillations lead to Wannier-Stark localisation of the dynamics [26, 38], with the resulting evolution displayed in Fig. 2.2. While it looks superficially similar to the case of confining quenches illustrated in Fig. 2.1, we reveal significant differences in the underlying mechanism behind the escaping fronts.
@@ 3.1 Confinement: mesons
It is clear from Fig. 3.1 that these functions have a maximum value inside the Brillouin zone. The existence of this maximum is guaranteed by the Lieb-Robinson theorem [14], and its value depends on the species $n$. For collisionless mesons (large $n$), the independence of their energy from their momentum $K$ implies that these mesons are stationary irrespective of the value of $K$, which is nicely explained by the semi-classical description. In addition, the velocities monotonously decrease with $n$; hence, the light-cone bounding the escaping fronts in Fig. 1(a) corresponds to the maximum velocity of the lightest meson species $M_{1}$.
@@ 3.2 Anti-confinement: bubbles
Note that in contrast to the mesons, the sign of the velocity is opposite to that of the momentum. However, these velocities still have finite maxima; the largest of these maximum velocities corresponds to the first bubble $B_{1}$, which is the one corresponding to the light-cone boundary of anti-confining quenches in Fig. 2(a).
@@ 4.1 Initial state for $h_{l}=0$
|$\ket{+}_{h_{t},h_{l}=0}=\frac{1}{\sqrt{2}}\left(|0\rangle_{\mathrm{NS}}+|0\rangle_{\mathrm{R}}\right)\,.$ (4.1)
@@
$|\Psi(0)\rangle=\hat{\sigma}^{z}_{L/2}\ket{+}_{h_{t},h_{l}=0}=\frac{1}{\sqrt{2}}\left(|\Psi(0)\rangle_{\mathrm{NS}}+|\Psi(0)\rangle_{\mathrm{R}}\right)\,,$ (4.2)
@@
$\hat{\sigma}^{z}_{L/2}\ket{-}_{h_{t},h_{l}=0}=\frac{1}{\sqrt{2}}\left(|\Psi(0)\rangle_{\mathrm{NS}}-|\Psi(0)\rangle_{\mathrm{R}}\right)\,,$ (4.7)
@@
|$\ket{-}_{h_{t},h_{l}=0}=\frac{1}{\sqrt{2}}\left(|0\rangle_{\mathrm{NS}}-|0\rangle_{\mathrm{R}}\right)\,.$ (4.8)
@@ 4.2 Matching the meson wave functions
In this formula, $\beta_{n}(K)_{L}$ is the expected, non-physical overall phase that needs to be fixed for every state separately and depends on accidental phase choices for the eigenvectors obtained using numerical diagonalisation. [^4] While the knowledge of these phases is not essential in our considerations, they must be taken into account if one wishes to reproduce the time evolutions displayed in Figs. 1(a) and 2(a) in terms of the overlaps. However, there is a non-trivial momentum-dependent relative phase factor in (4.10), which can be attributed to a relative phase between the Bogolyubov fermions $\eta_{q}$ definition on the chain and those in the quantum mechanical description. Denoting the latter by $a_{q}$, the matching rules (4.12) correspond to the relation
@@ 4.3 Meson overlaps
which has a small global component. An example of such a quench is presented in Fig. C.1 a) of Appendix C, for the parameters $h_{t}=0.25$ and $h_{l}=0.2$. To the naked eye, there is no obvious difference between the time evolutions of the longitudinal magnetisation shown in Fig. C.1 a) and of the purely local quench shown in the second panel of Fig. 1(b). Indeed, the deviation turns out to be suppressed by around $3$ orders of magnitude compared to the signal from the localised component of the purely local quench.
@@
$\mathcal{C}_{n}(K)=\frac{1}{\sqrt{L}}\sum\limits_{\begin{subarray}{c}k\,\in\mathrm{NS/R}\\ k\neq 0,\pm K/2\end{subarray}}\sigma_{K/2-k}\sigma_{K/2+k}{\psi}_{n,K}^{*}(k)_{L}$ ${\mathcal{K}}(K/2-k,K/2+k)_{L}$
@@
Figure 4.1: Comparison of the meson overlaps with the (flipped) true vacuum state $\sigma^{z}_{L/2}\ket{+}_{h_{t},h_{l}=0}$ calculated by applying exact diagonalisation (shown by crosses), and the same calculated via the semi-analytic method given by Eq. (4.17) (shown by continuous curves) for the cases $h_{t}=0.25$, $h_{l}=0.1$ and $h_{l}=0.3$.
@@
The first two terms describe zero-momentum mesons building up the localised oscillations visible in Fig. 2.1. At the same time, the sum contains single mesons in a superposition of a left-moving and a right-moving state with the same velocity contributing with equal weights, corresponding to ”Schrödinger kitten” states. These states are present because the system’s initial state is not translationally invariant, unlike after global quenches. Due to the symmetry of the local quenches applied to the system, the expectation value of the total momentum is still required to be zero, which is ensured by the relation Eq. (4.18) that is indeed satisfied by the calculated overlaps. The emergence of the prominent light cone is explained by these ”Schrödinger kitten” states escaping confinement since the initial state is decisively dominated by single meson states and the calculated overlaps of the ”Schrödinger kitten” states are of the same order of magnitude as the overlaps of the zero-momentum meson states building up the localised oscillations as shown in Fig. 4.1.
Figure 4.2: Overlaps of meson states |$\ket{M_{n}(K)}$ with the state $\hat{\sigma}^{z}_{L/2}\ket{+}_{h_{t},h_{l}=0}$ (given by (4.2)), calculated via the semi-analytic method given by Eq. (4.17), as a function of $n$ at different values of momentum $K$.
@@ 5 Escaping fronts in anti-confining local quenches
$\overline{\mathcal{C}}_{n}(K)$ $=\frac{1}{\sqrt{L}}\sum\limits_{\begin{subarray}{c}k\,\in\mathrm{NS}\end{subarray}}\sigma_{K/2-k}\sigma_{K/2+k}\overline{{\psi}}_{n,K}^{*}(k)_{L}{\mathcal{K}}(K/2-k,K/2+k)_{L}$
@@
Figure 5.1: Comparison of the bubble overlaps with the (flipped) false vacuum state of the pure transverse system $\sigma^{z}_{L/2}\ket{-}_{h_{t},h_{l}=0}$ calculated from exact diagonalisation (crosses) with the same calculated via the semi-analytic method given by Eq. (5.2) at the parameter values $h_{t}=0.25$ and $h_{l}=0.1$ (continuous curves).
@@
Figure 5.2: Overlaps of bubble states |$\ket{B_{n}(K)}$ with the with the state $\sigma^{z}_{L/2}\ket{-}_{h_{t},h_{l}=0}$ (given by (4.7)), calculated via the semi-analytic method given by Eq. (5.2), as a function of $n$ at different values of momentum $K$.
@@
The latter conclusion is further supported by the fact that the maximum velocity of the smallest bubble $B_{1}$ matches the escaping fronts well, as shown in Fig. 2(a).
To sum up, for the anti-confining quenches, dynamical confinement is replaced by Wannier-Stark localisation but again, it is the superpositions of left and right-moving states that escape it, with the only difference that these are bubble states instead of mesonic ones.
@@ 6 Conclusions and outlook
In contrast with global quenches, our simulations show that after local quenches are obtained by flipping a single spin in the ground state, the magnetisation has a significant propagating component besides the localised one. To determine the origin of these fronts, we computed the overlaps of meson states with the initial state using a semi-analytic approach, which works for any chain length, including the thermodynamic limit. The conclusion is that the light-cone fronts originate from states containing a single meson excitation in superpositions of left and right-moving mesons (mainly the lightest one) with equal weights, which is in contrast to global quenches where quasi-particle pairs trace out the light cones and are strongly suppressed by confinement. The interpretation of the global quench scenario is closer to our classical intuition since the two branches of the light cone are traced out by pairs of quasi-particles with opposite momentum. However, in local quenches, it is one-particle states in the superposition of left and right moving momenta with equal amplitudes that are responsible for the presence of the emergent light cone.